Use row reduction to determine whether the set of vectors {(1,2,0), (0,1,-1),(1,1,2)} is linearly independent in                                                                 


1
Expert's answer
2021-09-06T07:03:40-0400

"\left( {\begin{matrix}
1&2&0\\
0&1&{ - 1}\\
1&1&2
\end{matrix}} \right)\begin{matrix}
{}\\
{}\\
{ - I}
\end{matrix} \sim \left( {\begin{matrix}
1&2&0\\
0&1&{ - 1}\\
0&{ - 1}&2
\end{matrix}} \right)\begin{matrix}
{}\\
{}\\
{ + II}
\end{matrix} \sim \left( {\begin{matrix}
1&2&0\\
0&1&{ - 1}\\
0&0&1
\end{matrix}} \right)"

Since the rank of the resulting matrix is ​​3, the system of these vectors is linearly independent